OpenStudy (anonymous):

Simplify each expression: 1. (3x^2)(-5x^-1)^2 2. ( 5y^4/2y^6)^3

ganeshie8 (ganeshie8):

use below exponent/power properties : \((ab)^n = a^n * b^n\) \(a^m * a^n = a^{m+n}\)

OpenStudy (anonymous):

is it 75 for the first one

ganeshie8 (ganeshie8):

1. \((3x^2)(-5x^{-1})^2\) \((3x^2)(-5)^2 (x^{-1})^2\) \((3x^2) (25)(x^{-2})\) \(25 (3x^2) (x^{-2})\) \(75 (x^2) (x^{-2})\) \(75 x^{2-2})\) \(75 x^0\) \(75 \)

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

for the second one is it 125 y^30/8

OpenStudy (anonymous):

thanks so much

ganeshie8 (ganeshie8):

for the second one you should get : \(\large \dfrac{125}{8y^6}\)

OpenStudy (anonymous):

thanks